# Hopf formulae for homology of skew braces

### v1

@@ 5. The subvariety of abelian groups
Let $A$ be a skew brace, we denote by $A^{\prime}$ the subgroup of $A_{+}$ generated by the elements $a*b=-a+a\circ b-b$ and $a+b-a-b$ for all $a,b\in A$.
@@
Let $A$ be a skew brace, and let $I\subset A$ be an ideal of $A$. We define the relative commutator $[I,A]_{\mathsf{Ab}}$ to be the additive subgroup generated by the set
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal of $A$. Then, $[I,A]_{\mathsf{Ab}}$ is an ideal of $A$.
@@
Since $J=[I,A]_{\mathsf{Ab}}\subset I$, $J$ contains the commutator subgroup $[J_{+},A_{+}]$ so $J_{+}$ is normal in $A_{+}$. In addition, $J*A\subset J$. To see that $A*J\subset J$, it is enough to show that for $k\in J$ and $b\in A$, $b*k\in J$. Since $-b+b\circ k\in I$, in $A_{+}/J_{+}$ one has
@@
An extension $(A,f)$ is central with respect to the adjunction (5.1) if and only if $[\operatorname{Ker}(f),A]_{\mathsf{Ab}}=0$.
@@
Since $(A,f)$ is central, by definition $a*b=R(s)(a*b,0)=R(t)(a*b,0)=0$. The first part of the proof of Theorem 4.13, where it is shown that $[a,b]_{+}=0$, still applies here. The proof that $[a,b]_{\circ}=0$ is also similar, and is left to the reader.
@@
Similarly, one has that $[a,a_{1}]_{\circ}\circ[b,b_{1}]_{\circ}^{\prime}=0$, which implies that the restrictions $R(s)$ and $R(t)$ of the first and the second projections $s$ and $t$ (with the notations from diagram (2.2)) are equal, as desired. ∎
One can ask whether the commutator $[I,A]_{\mathsf{Ab}}$ defined above coincides with the Huq commutator of normal subobjects defined in [15]. We shall now see that this is indeed the case.
Given an ideal $I$ of a skew brace $A$, consider the set-theoretic map $c\colon I\times A\rightarrow A$ defined by $c(i,a)=i+a$ for any $i\in I$ and $a\in A$. The Huq commutator $[I,A]_{\mathsf{Huq}}$ is the smallest ideal $J$ of $A$ such that the composite map
(5.2) ${{I\times A}}$${A}$${{A/J}}$$\scriptstyle{c}$$\scriptstyle{\pi}$
@@
By definition of $[I,A]_{\mathsf{Ab}}$, the set-theoretic map
${{I\times A}}$${A}$${{A/[I,A]_{\textsf{ab}}}}$
@@
is a homomorphism of skew braces, so that $[I,A]_{\mathsf{Huq}}\subset[I,A]_{\mathsf{Ab}}.$ For the other inclusion, consider any quotient $\pi\colon A\rightarrow A/J$ having the property that the composite map $\pi c$ in (5.2) is a homomorphism of skew braces. For any $i\in I,a\in A$,
@@
hence $\overline{[i,a]_{\circ}}=\overline{0}$. Similarly, one checks that $\overline{[i,a]_{+}}=\overline{0}$, and $\overline{i*a}=\overline{0}$, so that all the generators appearing in the Definition 5.2 of $[I,A]_{\mathsf{Ab}}$ must be sent to $\overline{0}$ by $\pi$. It follows that $[I,A]_{\mathsf{Ab}}\subset J$, hence in particular $[I,A]_{\mathsf{Ab}}\subset[I,A]_{\mathsf{Huq}}$, as desired. ∎
@@ 6. Hopf formulae for homology
00{0}${K}$${F}$${B}$00{0}$\scriptstyle{f}$
of a skew brace $B$, where $F={\mathbb{G}}(B)$ is the “free” skew brace on the underlying set of $B$. Consider also the ideal $[F,F]_{\mathsf{Grp}}=F*F$ of $F$ as defined in Section 3 and the ideal
$[K,F]_{\mathsf{Grp}}=\langle\{a*b,\quad b*a,\quad c+b*a-c\,\mid\,a\in K,b\in A,c\in A\}\rangle_{F}.$
@@
${{\mathsf{H}_{2}(A,F)}}$${{\mathsf{H}_{2}(B,F)}}$${{\frac{K}{[K,A]_{\textsf{Grp}}}}}$${{\mathsf{H}_{1}(A,F)}}$${{\mathsf{H}_{1}(B,F)}}$${0.}$$\scriptstyle{\mathsf{H}_{2}(f)}$$\scriptstyle{\mathsf{H}_{1}(f)}$
The same kind of results can be stated by replacing $[F,F]_{\mathsf{Grp}}=F*F$ with $[F,F]_{\mathsf{RadRng}}$ and $[K,F]_{\mathsf{Grp}}$ with $[K,F]_{\mathsf{RadRng}}$, so that the second homology radical ring ${\mathsf{H}}_{2}(B,G)$ of $B$ (with coefficients in the reflector $G\colon\mathsf{SKB}\rightarrow\mathsf{RadRng}$) is given by
@@
with coefficient functor $\mathsf{br}\colon\mathsf{SKB}\rightarrow\mathsf{BR}$, where the relative commutators $[F,F]_{\mathsf{BR}}$ and $[K,F]_{\mathsf{BR}}$ are described in Remark 4.14.
@@
$\mathsf{H}_{1}(f)\colon\mathsf{H}_{1}(A,\mathsf{ab})\rightarrow\mathsf{H}_{1}(A,\mathsf{ab})$
@@
$\mathsf{H}_{2}(f)\colon\mathsf{H}_{2}(A,\mathsf{ab})\rightarrow\mathsf{H}_{2}(A,\mathsf{ab})$

### v2

@@ 5. The subvariety of abelian groups
Let $A$ be a skew brace, we denote by $A^{\prime}$ the subgroup of $A_{+}$ generated by the elements $a*b=-a+a\circ b-b$ and a+baba+b-a-b for all $a,b\in A$.
@@
Let $A$ be a skew brace, and let $I\subset A$ be an ideal of $A$. We define the relative commutator $I,A]_{\mathsf{Ab}}$ to be the additive subgroup generated by the set
@@
Let $A$ be a skew brace, and $I\subset A$ an ideal of $A$. Then, $I,A]_{\mathsf{Ab}}$ is an ideal of $A$.
@@
Since $I,A]_{\mathsf{Ab}}\subset I, JJ$ contains the commutator subgroup $J_{+},A_{+}]$ so $J_{+}$ is normal in $A_{+}$. In addition, $J*A\subset J$. To see that $A*J\subset J$, it is enough to show that for $k\in J$ and $b\in A$, $b*k\in J$. Since $b+b\circ k\in I$, in $A_{+}/J_{+}$ one has
@@
An extension $(A,f)$ is central with respect to the adjunction (5.1) if and only if $\operatorname{Ker}(f),A]_{\mathsf{Ab}}=0$.
@@
Since $(A,f)$ is central, by definition ab=R(s)(ab,0)=R(t)(ab,0)=0a*b=R(s)(a*b,0)=R(t)(a*b,0)=0. The first part of the proof of Theorem 4.13, where it is shown that $a,b]_{+}=0$, still applies here. The proof that $a,b]_{\circ}=0$ is also similar, and is left to the reader.
@@
Similarly, one has that $a,a_{1}]_{\circ}\circ[b,b_{1}]_{\circ}^{\prime}=0$, which implies that the restrictions $R(s)$ and $R(t)$ of the first and the second projections $s$ and $t$ (with the notations from diagram (2.2)) are equal, as desired. ∎
One can ask whether the commutator $I,A]_{\mathsf{Ab}}$ defined above coincides with the Huq commutator of normal subobjects defined in [15]. We shall now see that this is indeed the case.
Given an ideal $I$ of a skew brace $A$, consider the set-theoretic map $c\colon I\times A\rightarrow A$ defined by $c(i,a)=i+a$ for any $i\in I$ and $a\in A$. The Huq commutator $I,A]_{\mathsf{Huq}}$ is the smallest ideal $J$ of $A$ such that the composite map
[…]
@@
By definition of $I,A]_{\mathsf{Ab}}$, the set-theoretic map
[…]
@@
is a homomorphism of skew braces, so that $I,A]_{\mathsf{Huq}}\subset[I,A]_{\mathsf{Ab}}.$ For the other inclusion, consider any quotient $\pi\colon A\rightarrow A/J$ having the property that the composite map $\pi c$ in (5.2) is a homomorphism of skew braces. For any $i\in I,a\in A$,
@@
hence $\overline{[i,a]_{\circ}}=\overline{0}$. Similarly, one checks that $\overline{[i,a]_{+}}=\overline{0}$, and $\overline{i*a}=\overline{0}$, so that all the generators appearing in the Definition 5.2 of $I,A]_{\mathsf{Ab}}$ must be sent to $\overline{0}$ by $\pi$. It follows that $I,A]_{\mathsf{Ab}}\subset J$, hence in particular $I,A]_{\mathsf{Ab}}\subset[I,A]_{\mathsf{Huq}}$, as desired. ∎
@@ 6. Hopf formulae for homology
[…]
of a skew brace $B$, where $F={\mathbb{G}}(B)$ is the “free” skew brace on the underlying set of $B$. Consider also the ideal $F,F]_{\mathsf{Grp}}=F*F$ of $F$ as defined in Section 3 and the ideal
$K,F]_{\mathsf{Grp}}=\langle\{a*b,\quad b*a,\quad c+b*a-c\,\mid\,a\in K,b\in A,c\in A\}\rangle_{F}.$
@@
[…]
The same kind of results can be stated by replacing $F,F]_{\mathsf{Grp}}=F*F$ with $F,F]_{\mathsf{RadRng}}$ and $K,F]_{\mathsf{Grp}}$ with $K,F]_{\mathsf{RadRng}}$, so that the second homology radical ring ${\mathsf{H}}_{2}(B,G)$ of $B$ (with coefficients in the reflector $G\colon\mathsf{SKB}\rightarrow\mathsf{RadRng}$) is given by
@@
with coefficient functor $\mathsf{br}\colon\mathsf{SKB}\rightarrow\mathsf{BR}$, where the relative commutators $F,F]_{\mathsf{BR}}$ and $K,F]_{\mathsf{BR}}$ are described in Remark 4.14.
@@
$\mathsf{H}_{1}(f)\colon\mathsf{H}_{1}(A,\mathsf{ab})\rightarrow\mathsf{H}_{1}(B,\mathsf{ab})$
@@
$\mathsf{H}_{2}(f)\colon\mathsf{H}_{2}(A,\mathsf{ab})\rightarrow\mathsf{H}_{2}(B,\mathsf{ab})$